OpenStudy (anonymous):

Bridget has 33 dimes and quarters in her piggy bank. If the value of these coins is $6.00, how many dimes and quarters does she have? Set up and solve a system of equations to solve the problem. Show your work.

OpenStudy (mathmate):

Use D to stand for dimes, and Q for quarters. Each dime is worth 10 cents, and each quarter is worth 25 cents. The system is therefore D+Q = 33 (total 33 coins) 10D+25Q=600 (total value of the coins) Most of the time, this type of can be solved with only one equation, as follows: D=number of dimes, 33-D=number of quarters, then for the total value 10D+25(33-D)=600 and solve for D.

OpenStudy (mathmate):

Even better, this type of problems can be solved without algebra, as follows: If all 33 coins are dimes, the total is worth is 825 cents. We are short (in value) by 825-600=270 cents. For each dime changed for a quarter, we increase the value by 25-10=15 cents. Number of exchanges required = 270/15=18 So there are 18 quarters, (33-18)=15 dimes.

OpenStudy (mathmate):

*If all 33 coins are dimes, the total is worth is 330 cents. We are short (in value) by 600-330=270 cents.

OpenStudy (anonymous):

Thank you so much I completely understood that!

OpenStudy (mathmate):

You're welcome! :)